### v2

# Energy landscape interpretation of universal linearly increasing absorption with frequency

Abstract
Absorption of elastic waves in complex media often depends on frequency in a linear way for both longitudinal and shear waves. This universal property occurs in media such as rocks, unconsolidated sediments, and human tissue. Absorption is due to relaxation processes at atomic scale up to the sub-micron scale of biological materials, and we argue that these processes are thermally activated. Unusual for ultrasonics and seismics, we can therefore express absorption as an integral over an activation energy landscape weighted by an energy distribution. The universal power-law property corresponds to a flat activation energy landscape corresponding to maximal randomness.
I Introduction
The amplitude of elastic waves undergoes attenuation which often follows a power-law in frequency, $\omega$:
$|u(x,\omega)|\propto\mathrm{e}^{-\alpha(\omega)x},\enspace\alpha(\omega)=\alpha_{0}\omega^{y},$ (1)
where ${u}$ is particle velocity and $x$ is distance. The two mechanisms for attenuation, $\alpha$, are viscoelastic absorption leading to heating, and scattering of energy. This paper is only concerned with mechanisms for absorption, although medium properties may also be inferred from the scattering, [1]. In complex media absorption often increases linearly with frequency, i.e. $y=1$, [2, Sect. 5.1], as the many examples that follow demonstrate.
In seismology, many nuclear-explosion and earthquake data sets in the range $10^{-3}$ to 10 Hz appear to have a constant quality factor, $Q$, for shear waves [3]. Inverse Q, also called specific attenuation or internal friction, is defined as $Q^{-1}=2\alpha(\omega)c(\omega)/\omega$, [2, Sect. 2.3], and if dispersion can be neglected, $c(\omega)=c_{0}$, it follows that $Q^{-1}\propto\omega^{y-1}$. Therefore, constant-Q behavior is the same as linearly increasing absorption.
Constant-Q behavior is especially apparent after correction for bias due to an additive constant in the expression for $\alpha(\omega)$, which may be due to geometrical spreading, (de-)focusing, or scattering, [3]. Similar results are reported for seismic reflection data from vertical wells [4] and constant-Q behavior up to about 100 Hz is also common in seismic survey data [5].
It is also recognized that $Q$ is nearly frequency-independent over one to two decades of frequency in many solids, [6]. This property was noted for metals and nonmetals, for both compressional and shear waves, and for frequencies in the Hz, kHz, and MHz range.
Unconsolidated sub-bottom sediments represent a very different medium, but the same linearly increasing absorption of compressional waves is observed above about 2 kHz [7]. Below that frequency the exponent is closer to two. Even in this case the value for the exponent, $y$, will depend on whether shear-wave mode conversion in the form of a constant term in the expression for $\alpha(\omega)$ is compensated for, [8].
One of the best studied fields is compressional waves in the MHz range in medical ultrasound, [9]. A recent review says that an “early consensus emerged that a power law fit near 1 was adequate for absorption models of soft tissues” [10]. As in the case of seismology, care is needed in interpreting measurements as defocusing due to phase aberrations may lead to an overestimation of absorption if not compensated for [11].
The medical ultrasound field has provided insight into the spatial scale where absorption takes place. Grinding liver tissue hardly changes the absorption, so it is apparent that the mechanisms operate ”on a level of organization smaller than that defined by cells, cell nuclei, and mitochondria”, i.e., size less than about 3 microns, [12]. Absorption in canned evaporated milk also follows a linear frequency law [13], indicating that casein micelles at the sub-micron size level play an important role. Similarly, absorption of compressional waves in blood in the 0.8-3 MHz range was due to the presence of proteins [14].
A related phenomenon is how cells or tissue react to shear excitation. The response often varies with $\omega^{\beta}$, where $\beta$ is in the range 0.1 to 0.5, [15], [16]. According to results cited in the Supplement, this corresponds to an absorption that varies with $y=\beta+1$ and therefore $y$ in this case will fall in the range 1.1 to 1.5.
The property that the exponent $y$ is near unity over a wide range of materials is nearly universal, both for compressional and shear waves, as noted in e.g. [17]. The exponent may be larger than unity, but never above 2, the viscous case, and rarely below 1. The temperature range of interest is approximately 0 to 40 C.
This paper is concerned with viscoelastic absorption, but there are other mechanisms which will not be discussed such as friction in cracks in rocks, fluid flow in micropores in consolidated sediments or bone, squirt flow in unconsolidated sediments, and nonlinearity.
An early study of absorption in erythrocytes suggested that “chemical or structural relaxation processes are probably responsible for the attenuation” [18]. It has also been proposed that in polymer-like materials, thermal energy causes continuous change in the “interactions between macromolecules” [17]. In cell rheology, the theory of soft glassy materials has been shown to be applicable, [15].
We also notice that chemical and structural relaxation in e.g., seawater are thermally activated processes. Thermal activation over potential barriers of structural defects also plays an important role in e.g. glasses and rocks. This gives the motivation to transform the multiple relaxation formulations from the usual frequency and time domains into the energy domain. This will provide new insights for the case of elastic wave propagation in acoustics and seismics. Surprisingly, a simple uniform distribution of activation energies corresponds to the universal linearly increasing absorption.
II Background: Multiple relaxation
A relaxation process is characterized by a relaxation frequency, $\Omega$, or a relaxation time $\tau=1/\Omega$, and an absorption given by: ${\Omega\,\omega^{2}}/({\omega^{2}+\Omega^{2}})$. Attenuation increases with $\omega^{2}$ well below the relaxation frequency and is constant well above it. This expression can for instance be found from structural relaxation, [19], and from chemical relaxation, [20].
In a complex medium there are many elementary relaxation processes over a large spread of relaxation frequencies and absorption is:
$\alpha(\omega)=A_{0}\;\omega^{2}\int_{0}^{\infty}\frac{g_{\Omega}(\Omega)\,\Omega}{\omega^{2}+\Omega^{2}}\,\mathrm{d}\Omega,$ (2)
where $A_{0}$ is a constant. The weighting $g_{\Omega}(\Omega)$ in the relaxation integral has the form of a probability distribution function. The particular distribution given by $g_{\Omega}(\Omega)=\Omega^{y-2}$ will result in the power-law absorption of (1), [21, Sect. 3.241.2], [22]. This formulation does, however, not provide much insight to motivate why $g_{\Omega}(\Omega)$ should follow this particular power-law relation.
The integral of (2) can be transformed to be over relaxation times, $\tau$, by letting $g_{\Omega}(\Omega)=g_{\tau}(\tau)\;|\mathrm{d}\tau/\mathrm{d}\Omega|$:
$\alpha(\omega)=A_{0}\;\omega^{2}\int_{0}^{\infty}\frac{g_{\tau}(\tau)\;\tau}{1+\omega^{2}\tau^{2}}\mathrm{d}\tau,$ (3)
where the particular distribution $g_{\tau}(\tau)=\tau^{-y}$ will lead to the desired power-law absorption, [23]. This formulation may be easier to interpret as relaxation times may be related to length scales, $l=c\tau$, by means of the speed of propagation, $c$. Cell biomechanics may offer some insights, as power-law behavior of the shear modulus over about five decades of frequency is well documented. However, there seems not to be enough cell components in a hierarchy from the largest to the smallest to account for the observed power-law behavior, [24, 15, 25]. There is therefore limited physical insight to gain even from the formulation of (3).
III Thermally activated relaxation
In acoustics one is usually content with the descriptions of (2) and (3). The limitation, as noted, is that there is little insight to gain into why the relaxation processes are ”organized” to give power-law characteristics.
A closer look at the common mechanisms for intrinsic absorption is therefore warranted. In acoustics, they are viscosity, molecular thermal relaxation, heat conduction in monatomic gases, structural relaxation, and chemical relaxation, [26, Chap. 8], [2, Sect. 4.1]. Only the two last ones are relevant to properties of fluids and solids, our main interest in this paper. We therefore start by reviewing the well-established model for seawater, as an example of a medium where structural (also called segmental) relaxation and chemical relaxation dominate.
III.1 The Arrhenius law
Absorption in seawater has three main components. Two of them are due to B(OH)3 and MgSO4 which both contribute two-state chemical equilibrium reactions. Their relaxation frequencies are in the kHz and tens of kHz range respectively. Relaxation time in these reactions is related to activation energy, $E_{a}$, according to the Arrhenius relation, [20]:
$\tau=\tau_{0}\mathrm{e}^{E_{a}/k_{B}T},$ (4)
where $T$ is absolute temperature, $k_{B}$ is Boltzmann’s constant, and $\tau_{0}$ characterizes the smallest time scale probed. It should be noted that in underwater acoustics, chemical relaxation is usually parameterized with temperature in Celcius rather than Kelvin, [27], obscuring the fact that these processes are thermally activated.
The most important contribution to absorption in water is from structural relaxation of H2O molecules. In distilled water a broken-down structure of clusters is dynamically changed by an incoming sound wave and relaxation takes place as clusters of different sizes interact. Based on a two-state energy model for H2O molecules, the Arrhenius law describes the transition rates, [19]. The relaxation frequency is in the THz range.
The Arrhenius law point to an activation energy landscape interpretation of chemical and structural relaxation taking place at the molecular or molecular cluster level, i.e. at the nanometer scale. For now, that only covers some of the cases mentioned in the Introduction, but let us anyway pursue the consequences of this view point.
III.2 Transformation of relaxation integral
The validity of the Arrhenius relation is a key insight that allows us to transform the previous relaxation integrals. Equation (3) can be transformed by using (4) in combination with $g_{\tau}(\tau)=g_{E}(E_{a})/k_{B}T\cdot\;|\mathrm{d}E_{a}/\mathrm{d}\tau|$:
$\alpha(\omega)=\frac{A_{0}}{k_{B}T}\,\omega^{2}\int_{0}^{\infty}\frac{g_{E}(E_{a})\tau(E_{a})}{1+\omega^{2}\tau^{2}(E_{a})}\mathrm{d}E_{a},$ (5)
where $g_{E}(E_{a})$ is an energy distribution function. It should be noted that activation energy and its distribution function are physical parameters that lend themselves to interpretation much more than distributions of relaxation frequencies or relaxation times.
III.3 Glass
Equation (5) is common for describing the effect of structural defects in glassy media, [28] with relaxation time given by the Arrhenius relation, [29]. A glass can be considered to be frozen in an energy basin with many local minima. The height of the barriers between them, and the energy difference between the minima will determine properties, [30]. Structural relaxation takes place due to perturbations of the landscape and this is the main cause of absorption. We are mainly concerned with normal temperatures (about 270 - 310 K), and then tunneling [31] can be neglected and the classical model with the thermal activation rate of (4) describes the relaxation.
This describes what happens in structural glasses, but is also applicable to rocks [32], where most minerals consist of crystalline grains with amorphous grain boundaries with structure similar to glasses.
III.4 Energy distribution
The Arrhenius relation allows the transformation between the frequency, time and energy relaxation integrals, (2), (3), and (5). The energy distribution may be written as:
$g_{E}(E_{a})={\Omega}\;g_{\Omega}(\Omega)={\tau}\;g_{\tau}(\tau).$ (6)
In the case of the power-law absorption of interest here, we have
$g_{E}(E_{a})={\Omega^{y-1}}.$ (7)
The most interesting case is the energy distribution corresponding to linearly increasing absorption, $y=1$. Surprisingly, that corresponds to a flat activation energy distribution.
A general distribution can be further investigated by combining (7) with (4), giving:
$g_{E}(E_{a})={\Omega_{0}^{y-1}}\mathrm{e}^{\frac{-E_{a}}{k_{B}T}(y-1)}.$ (8)
This result is plotted in Fig. 1. The most apparent feature is the flat distribution for $y=1$, as noted. In addition the case for $y>1$ is also interesting. It corresponds to an exponential energy distribution where lower energies are given more weight than higher energies.
Figure 1: Normalized energy distribution function of (8) plotted for power-law exponents $y$ in the range from 0.9 to 1.3 and for $\tau_{0}^{-1}=2\pi\cdot 10^{9}$.
So far we have given arguments for why (8) is valid for media where relaxation takes place at the nanometer scale. In the next section the size scale will be expanded.
IV Energy landscapes and biological materials
Interestingly, the energy landscape interpretation is used over a much larger range of scales in [33, Chap. 1]. Wales’s book starts by outlining three different fields: “The structure and dynamics of atomic and molecular clusters, the folding of proteins, and the complicated phenomenology of glasses are all manifestations of the underlying potential energy surface”
The first and last of these fields have already been mentioned as structural and chemical relaxation and as the effect of structural defects in glasses. Surprisingly, energy landscapes are used even for proteins and folded proteins, including micelle formation [34]. That would include most, if not all, of the materials mentioned in the Introduction, from the nanometer scale up to the sub-micron range.
IV.1 Validity of the Arrhenius relation
One thing is that properties are described by an energy landscape, another is the validity of the Arrhenius relation. As stated in [35], the Arrhenius relation comes from statistical mechanics and is valid for a system which transitions from one metastable state to another. They argue that although these assumptions are not necessarily valid in biological tissue, analogs to the relevant parameters “exist in cells and likely govern cell motility.”
Likewise, [36] argues that proteins also can be regarded as two-state systems. They cite as examples “an ion channel can be open or closed, a hemoglobin or myoglobin protein can have bound oxygen or not.” The Arrhenius law is therefore used in these fields as well, although the physical basis is not as solid as for processes at the atomic and molecular scale.
IV.2 Soft glassy materials and glass
There is independent evidence that human tissue cells under the influence of shear may be modeled with soft glassy rheology [37], [15], [38]. This lends support to the just mentioned descriptions of biological materials,
The properties of soft glasses correspond to those of glasses at temperatures between the glass temperature and the melting point, $T_{g}<T<T_{m}$. The cell’s mechanical properties are determined by the crowded interior of the cell. This is analogous to what takes place in a colloidal suspension and leads to the complex shear modulus following a weak power law over several frequency decades with a near constant power law exponent, similar to that of Supplement Eq. (S7). Cells are examples of a disordered metastable material which exists in a state far from thermodynamic equilibrium. The energy landscape is comprised of the binding energy between neighboring proteins. A deformation due to an incoming wave may cause a hop between energy wells where the deformation energy is taken from the wave, i.e. leading to heating and absorption of wave energy.
IV.3 Sensitivity to energy distribution in the band-limited case
As is clear from the examples of the Introduction, the power-laws are only observed over a limited bandwidth. In [39] and [23] it is shown that (2) and (3) result in a good fit to (1) even in that case. This means that each component in the integral mainly affects frequencies in the vicinity of its relaxation frequency. Band limiting to a range from $\Omega_{L}$ to $\Omega_{H}$ in (2), implies that the asymptotes of the power-law absorption will be
$\alpha(\omega)\propto\begin{cases}\omega^{2},&\omega\ll\Omega_{L},\\ \omega^{y},&\Omega_{L}\ll\omega\ll\Omega_{H},\\ \alpha_{\infty},&\Omega_{H}\ll\omega.\end{cases}$ (9)
A high-frequency limit like that of (9) may in fact be required for physical reasons as a passive medium requires that the absorption should not increase faster than $\omega^{1}$ as $\omega$ approaches infinity, [40]. As noted, the low-frequency limit is relevant for modeling e.g., sub-bottom sediments, [41]. The exponential relationship of the Arrhenius law also means that a band-limiting corresponds to a relatively narrow range of activation energies.
In the band-limited case, the skewed distribution of (8) may be approximated by the first term in a Taylor series about a point in the middle of the energy range, $E_{c}$. This point corresponds to a frequency $\Omega_{c}=\sqrt{\Omega_{L}\Omega_{H}}$ and the linear approximation of the distribution is:
$g_{E}(E_{a})\approx g_{E}(E_{c})\left[1+\frac{1-y}{T}\frac{E_{a}-E_{c}}{k_{B}}\right],$ (10)
which is an acceptable approximation for a narrow frequency range and a power-law exponent near unity. Fig. 2 shows an example that demonstrates that when (1) is only given over two decades, the exact shape of the energy distribution is not critical. The difference between the absorption for the exact and the approximated linear case, are minor and most likely often smaller than the measurement error. The figure also illustrates the lower and upper asymptotic values given by (9).
Figure 2: Band-limited example with power-law $y=1.3$. Comparison between exact energy distribution (blue, solid line), and a linear approximation (black, dashed line), and with reference curves with slopes $y=1.3$ and $y=2$ above them.
The inset shows the exact (blue, solid line) and linearized (black, dashed line) energy distributions with $\tau_{0}^{-1}=2\pi\cdot 10^{8}$, $T=300$ K, $f_{L}=10^{5}$ and $f_{H}=10^{7}$ Hz corresponding to $E_{a}/k_{b}$ between 2072 and 691 K.
V Discussion
Glasses have a universal property at low temperatures, 0.1 to 10 K, where $Q^{-1}(\omega;T)$ is found to be nearly independent of temperature T as well as frequency $\omega$, [42]. The property that elastic wave absorption depends on frequency in a linear way around room temperature is a similar universal property which results from a flat activation energy distribution.
An alternative way of deriving an energy distribution similar to (8) is found in the theory for soft glassy materials, [43], [44] where the medium was modeled as a Maxwell-Wiechert model as in Fig. (S1) and Eq. (S5) of the Supplement.
That theory also gives a material description in the form of a partial differential equation which expresses how regions rearrange to new positions, valid for low frequencies. The main variable is $g_{E}(E_{a})$, the probability for finding an element trapped in a barrier between the two wells of height $E_{a}$.
The theory contains a constant which is an attempt frequency, and an activation factor on the same form as the Arrhenius equation. The theory is expressed in normalized units with a central parameter being the mean-field noise temperature, $y$. A glass transition occurs at $y=y_{g}=1$ and the material approaches the fluid state for $y=2$. Another input is a prior distribution of traps which it is argued has an exponential tail, $p_{a}(E_{a})=\exp(-E_{a}/y_{g})$, [45].
The equilibrium distribution of energies above the glass transition is given by $g_{E}(E_{a})\propto\mathrm{e}^{E_{a}/y}p_{a}(E_{a})$, and although the end result is not stated explicitly in [44, Sect. IV.A], it is an exponential distribution:
$g_{E}(E_{a})\propto\mathrm{e}^{\frac{-E_{a}}{y}(y-1)}.$ (11)
Since $y$ corresponds to $k_{b}T$, this expression is analogous to (8). In addition, it states that the parameter $y$ changes with temperature. New results are now appearing with ultrasound properties of tissue at low temperature that potentially could confirm this, [46].
Further it is demonstrated how this model leads to a dynamic modulus ${E}(\omega)={E}^{\prime}(\omega)+\mathrm{i}{E}^{\prime\prime}(\omega)$ where both the real and the imaginary components are proportional to $\omega^{y-1}$, as for the fractional Kelvin-Voigt model of Supplement Fig. (S2) and Eq. (S7).
Thus the noise temperature and the fractional order have a simple relationship, $y=\beta+1$. The soft glassy model therefore provides an interpretation of the fractional order, $\beta$ as well as for $y$. The special case of concern in this article is found in the limit, as $y$ approaches one, i.e. the material approaches the glass temperature. It is also evident that the soft glassy model result of (11) for the energy distribution resembles our result (8) which we found in an independent way.
As mentioned in the Introduction, the typical range for $y$ for shear waves in biological tissue is up to $1.5$. This corresponds to an energy distribution which is skewed towards lower energies, see Fig. 1. One explanation for why it is not flat could be that the Arrhenius relation is not completely valid for some biological materials. Variations such as the stretched exponential Arrhenius relation could be relevant for further investigation.
VI Conclusion
It is remarkable that elastic wave absorption depends on frequency in a linear way around room temperature universally across applications as diverse as seismology, seismics, subbottom acoustics, and medical ultrasound.
Such absorption is the result of a large number of relaxation processes, expressed by a weighted relaxation integral over frequency or over time. It is however hard to argue physically why the particular weighting that gives rise to power-law absorption, should occur. We have argued here, based on properties of atomic and molecular clusters, proteins, and glasses that an energy landscape formulation is fundamental. Further the Arrhenius expression for activation energy, despite being formally derived in statistical mechanics, is valid. This enables the transformation of the multiple relaxation formulation to an integral over energies in an energy landscape. The formulation derived here is not common in e.g. ultrasonics and seismics.
A macroscopic property, absorption, is linked to properties at the mesoscale level, the shape of the energy landscape. The interesting case of linearly increasing absorption corresponds to a flat activation energy distribution, i.e., all energies are equally probable. A flat energy distribution indicates a form of equilibrium, and it is not unlikely that properties of the energy landscape will enable a deeper understanding of both the conditions for linearly increasing absorption with frequency as well as the origin of power-law relaxation responses in general.
VII Acknowledgments
We want to thank Svein-Erik Måsøy and Kevin Parker for valuable discussions of the topic of this paper, and to Yuri Galperin, Sven Peter Näsholm and Ralph Sinkus for comments on an early version of the manuscript.